OpenStudy (anonymous):

Help? If f(x) = |x – 5| + 2, find f(3)

OpenStudy (anonymous):

if you go to mathway.com and type in your question it will give you the answer. i use it all the time. :)

OpenStudy (anonymous):

Found the answer myself~! its 4 c:

OpenStudy (anonymous):

Just plug in 3 into x. 3^(2)/(3+3) 9/6 3/1 3 Or u can us this f(3) = 3^2/(3+3) = 9/6 = 3/2

OpenStudy (anonymous):

that is good:)

OpenStudy (anonymous):

Just plug in 3 into x. 3^(2)/(3+3) 9/6 3/1 3